OpenStudy (calculator):

Integrate 1/(sqrt(1-x^2)) Explain please :D

OpenStudy (callisto):

\[\int\limits (\sqrt{1-x ^{2}} )^{-1}dx = (1/x )(\sqrt{1-x ^{2}})+c\]

OpenStudy (callisto):

i'm not sure for that..

OpenStudy (anonymous):

look in your book and you will find a function whose derivative is \[\frac{1}{\sqrt{1-x^2}}\] hint, it is section on derivatives of inverse trig functions

OpenStudy (calculator):

my book did not say anything about substituting inverse trigonometry

OpenStudy (anonymous):

not a substitution

OpenStudy (anonymous):

\[\frac{d}{dx}[\sin^{-1}(x)]=\frac{1}{\sqrt{1-x^2}}\] and so \[\int\frac{1}{\sqrt{1-x^2}}dx =\sin^{-1}(x)+C\]

OpenStudy (calculator):

is it a formula?

OpenStudy (anonymous):

i am not sure i would call it a formula. it is a fact though

OpenStudy (agreene):

it's a definition.

OpenStudy (calculator):

because i dont understand how the trigonometry get into this equation

OpenStudy (anonymous):

if you take the derivative of any inverse function, you can do it via \[\frac{d}{dx}[(f^{-1})'(x)]=\frac{1}{f'(f^{-1}(x))}\]

OpenStudy (callisto):

Put \[x=\sin \theta\] dx = cosθ dθ \[\int\limits (\sqrt{1-x ^{2}})^{-1} dx = \int\limits (\sqrt{1-(\sin \theta)^2} )^{-1} \cos \theta d \theta = \int\limits d \theta = \theta+C =\sin^{-1} x+C\]

OpenStudy (callisto):

by trigo substitution sorry

OpenStudy (agreene):

\[\int \frac{1}{\sqrt{1-x^2}}dx ≡ \sin^{-1}(x)+C\]

OpenStudy (anonymous):

therefore \[\frac{d}{dx}[\sin^{-1}(x)]=\frac{1}{\cos(\sin^{-1}(x))}=\frac{1}{\sqrt{1-x^2}}\]

OpenStudy (anonymous):

it works for sure, but i would not use a trig sub for this because you are going around in circles. the derivative of arcsine is \[\frac{1}{\sqrt{1-x^2}}\] so the anti derivative of \[\frac{1}{\sqrt{1-x^2}}\] is arcsine they say the same thing

OpenStudy (calculator):

thanks @satellite73 and @Callisto !!!!!! i understand it now!!

OpenStudy (anonymous):

yw

OpenStudy (callisto):

sorry, can you explain why \[d/dx [\sin^{-1} x] = 1/\cos(\sin^{-1} x) \] i don't understand this :(

OpenStudy (anonymous):

sure

OpenStudy (agreene):

\[\frac{d}{dx}[(f^{-1})'(x)]=\frac{1}{f'(f^{-1}(x))}\]

OpenStudy (anonymous):

\[f\circ f^{-1}(x)=x\] \[f'(f^{-1}(x))(f^{-1}(x))'=1\] by the chain rule, therefore \[(f^{-1}(x))'=\frac{1}{f'(f^{-1}(x))}\] by algebra

OpenStudy (anonymous):

gives a method for finding derivatives of inveres function, reciprocal of derivative of original function evaluated at the inverse

OpenStudy (anonymous):

is one way to show that \[\frac{d}{dx}\ln(x) =\frac{1}{x}\]

OpenStudy (anonymous):

now you have \[f(x)=\sin(x), f'(x) = \cos(x), f^{-1}(x)=\sin^{-1}(x)\] so \[\frac{d}{dx}\sin^{-1}(x)=\frac{1}{\cos(\sin^{-1}(x))}\]

OpenStudy (callisto):

sorry to bother, but i don't understand the first 2 steps..

OpenStudy (anonymous):

\[f\circ f^{-1}(x)=x\]

OpenStudy (anonymous):

that one?

OpenStudy (callisto):

yup and the following step

OpenStudy (anonymous):

that is true by definition of an inverse function when you compose a function with its inverse you get the identity

OpenStudy (anonymous):

\[f\circ f^{-1}(x)=x\] by what it means to be an inverse

OpenStudy (anonymous):

\[e^{\ln(x)}=x\] \[\sin(\sin^{-1}(x))=x\] etc

OpenStudy (anonymous):

\[f'(f^{-1}(x))(f^{-1}(x))'=1\] is true because the derivative of the right hand side is clearly 1, and the derivative of the left hand side you get by the chain rule

OpenStudy (callisto):

thanks i understand this part now, but still have the other part, how to get \[[\cos (\sin^{-1} x)]^{-1} = (\sqrt{1-x ^{2}})^{-1}\]